# A Survey on Data Selection for LLM Instruction Tuning

### v2

@@ 3.4. Methods based on Small Models
Du et al. (2023) introduces the MoDS method, focusing on instruction selection through three criteria: quality (instructional data fidelity), coverage (variety of instruction types), and necessity ( impact of instruction on LLM fine-tuning). The process unfolds in four key steps. Firstly, employing a reward model to evaluate the instruction dataset quality, selecting a subset denoted as $D_{h}$, comprising instructions that surpass a predefined quality threshold. Secondly, utilizing the k-center-greedy algorithm (Sener and Savarese, 2018a) to identify seed instructions, which ensures a diverse and representative sample of the instruction dataset. Thirdly, fine-tuning a pre-trained LLM with the seed instructions. Subsequently, this refined model is applied to $D_{h}$ to generate a new dataset, $D_{inference}$. This dataset is then evaluated using the reward model, which identifies instructions critical for learning of the LLM by focusing on those with lower scores. A threshold is established to select Augmented Instruction Data, which is specifically tailored to enhance performance of the model.Finally, the seed instructions with the Augmented Instruction Data to form a high-quality instruction subset that is poised to facilitate effective fine-tuning of the LLM.
Chen et al. (2023b) proposes a coreset-based and task-related data selection method: firstly, obtain sentence embeddings of samples by pre-trained language models (like BERT), then select the center point by applied unsupervised clustering on these embeddings, and finally use the KCenterGreedy (Sener and Savarese, 2018b) algorithm to retrieve core samples from the given dataset. This method effectively reduces the required training data volume, maintaining or potentially enhancing model performance.
@@ 3.5. Comparative Analysis of Data Selection Methods
Model Dependence Model-agnostic and heuristic Requires supervised model training Depends on external black-box APIs (e.g., GPT-4) Uses small pre- trained models (e.g., BERT, TinyLM)
@@
Domain Adaptability Low High Medium Low to Medium
@@
Selection Effectiveness Variable High High Medium
@@
In terms of scalability, indicator-based methods have a clear advantage. They are generally fast, parallelizable, and well-suited for large-scale filtering tasks, as they do not require model inference. Trainable methods, by contrast, face serious scalability challenges due to the high computational cost of model training and inference. Black-box LLM-based methods suffer from scalability bottlenecks due to API latency, usage limits, and costs, making them impractical for processing massive datasets unless sufficient resources are available. Small model methods offer a middle ground—they are not as scalable as pure heuristics but can be applied to larger datasets with reasonable efficiency.
@@
When it comes to interpretability, heuristic indicator methods clearly lead, offering intuitive, transparent criteria for data selection (e.g., length, diversity, perplexity). This is important for tasks requiring human oversight or explainable decision-making. In contrast, both trainable and black-box LLM-based methods operate as opaque systems, making it difficult to trace why certain instances are selected. Small models offer moderate interpretability—they can be analyzed via model outputs and saliency techniques but are not as transparent as heuristics.
@@ 4.1. Wining Rate
To evaluate the effectiveness of the dataset selection methods, (Chen et al., 2024) define the winning rate of the LLM fine-tuned with less samples (i.e., LLM-sub) compared to the LLM fine-tuned with full samples (i.e., base LLM) and is calculated as:
(10) Win $\text{Win Rate}=\frac{Num(win)-Num(lose)}{Num(all)}+1,$
@@
LLM-sub represents the LLM fine-tuned on the subset that filtered from the training set by the selection methods while the base LLM often involves 2 types: i) fine-tuned on the full training set and ii) fine-tuned on the same-scale subset filtered by the regular selections (e.g. random sampling and instruction length). The outputs of LLM-sub and the base LLM are rated by the judge on a scale from 1 to 10, and usually GPT4 is employed as the judge. To address the positional bias of the judge, Li et al. (2024c) send the outputs of these two LLMs to the judge twice with different orders. According to Li et al. (2024c), the wining cases refer to LLM-sub outperforming the base LLM in both times or winning in one and tying in the other. The losing cases refer to LLM-sub lagging behind the base LLM in both times or tying in one and losing in the other. The wining rates of different selection methods on the testing benchmarks are summarized in Table 2.
Table 2. Performances of different selection methods on wining rate.
Selection methods Model pairs for comparison Training set Benchmark for testing (WS) Vicuna/Koala/WizardLM/Self-inst/LIMA total WS IFD (Li et al., 2024c) Llama-7B(5%), Llama-7B(full) alpaca 1.125/0.97/1.077/1/1.1 1.04 Llama-7B(10%), Llama-7B(full) 1.037/1.055/1.114/1.123/1.103 1.097 Llama-7B(15%), Llama-7B(full) 1/1.038/1.114/1.027/1.09 1.064 Llama-7B(10%), Llama-7B(full) WizardLM 1.1625/1.1278/1.1147/1.0278/1.1067 1.0971 Llama2-7B(5%), Llama2-7B(full) alpaca 1.5875/1.4889/1.4266/1.2937/1.4733 1.4311 Random Sampling Llama-7B(5%), Llama-7B(full) alpaca - 0.9 InstructionGPT4 (Wei et al., 2023) miniGPT4(6%), miniGPT4(full) cc_sbu_align - 1.167 Alpagsaus (Chen et al., 2024) Llama-7B(9k), Llama-7B(full) alpaca 1.2125/1.0222/1.0596/1.0556/- 1.0658 Llama-7B(9k), Llama-7B(3k) 1.1/1.183/1.082/1.17/- 1.074 Llama-7B(9k), Llama-7B(6k) 1.05/1.072/1.05/1.087/- 1.082 Llama-13B(9k), Llama-13B(full) 1.2125/1.0167/1.133/1.0198/- 1.074 MoDS (Du et al., 2023) Llama2-7B(2k), Llama2-7B(full) alpaca 1.7125/1.5111/1.4725/1.369/1.4933 1.4786 InstructionMining (Cao et al., 2023) Llama-7B(2k), Llama-7B(full) dolly - 1.088
  • ’-’ denotes values not reported in the original paper, and model(x) denotes tuning the model on the given training set using either x samples or x percentage of samples.
@@
Selection methods Training set(samples) Base model Benchmarks for testing total WS activeIT (Kung et al., 2023) selfinstruct(2k) Llama-7B 1.107 selfinstruct(full) 1.293 BBH/DROP/MMLU/Human-Eval/Avg DQ (Zhou et al., 2023b) alpaca(20%) Llama-7B 32.7/26.7/39.8/9.2/27.1 alpaca(full) 32.9/26.3/41.6/10/27.7 Vicuna RS/Vicuna WTR/Koala RS/ Koala WTR DIVERSEEVOL (Wu et al., 2023) Dolly(1k) Llama-7B 79.69/20/62.29/6.67 Dolly(full) 73.84/5/57.9/3.33 SelfIns(1k) 79.16/7.5/66.95/6.11 SelfIns(full) 73.03/2.5/69.5/3.89 HellaSwag/ARC/TruthfulQA/MMLU LIFT (Xu et al., 2023b) Mistral-7B 0.823/0.602/0.426/0.627 Platypus(15k random) 0.82/0.607/0.438/0.625 Platypus(15k) 0.844/0.643/0.49/0.645 RTE/CB/ANLI R1/ANLI R2/ANLI R3 coreset (Chen et al., 2023b) P3(0.5) Galactica-1.3B 74.73/73.21/49.6/41.9/43.75 P3(full) 76.17/75/44/35.7/39.42
  • dataset(x) denotes tuning the base model on the x samples of the given set. ✗denotes either the absence of a training set or a base model.
@@ 4.2. Inner Comparison
Methods Training set(samples) Base model MT-bench AlpacaEval gpt-4 8.99 95.28 gpt-3.5-turbo 7.94 91.36 alpaca-13B alpaca Llama-13B 4.53 - NUGGETS(Li et al., 2024a) alpaca(7.5k) Llama-7B 5.34 - TAGLM-13B-v1.0 (Lu et al., 2024) mixture(6k) Llama-13B 6.44±0.04 72.8 TAGLM-13B-v2.0 (Lu et al., 2024) Llama2-13B 6.55±0.02 - Instruction Length Llama-13B 5.89 - Random Sampling 5.84 - IFD(Li et al., 2024c) 5.91 - Instruction Node(Zhao et al., 2023) 5.65 - DEITA (Liu et al., 2024) mixture(10k) Llama2-13B 6.79 81.09
@@ 4.3. External Comparison
External comparison helps reveal the generalization ability of data selection methods across different base models and architectures. For example, as shown in Table 4, IFD(Li et al., 2024c) demonstrates robust performance when transferring from Llama-7B to Llama2-7B, indicating that the selected data subsets contain general instruction information beneficial to multiple models. However, this evaluation also exposes some limitations: certain methods that perform well in inner comparison show diminished advantages in external comparison, possibly due to overfitting to the training distribution of specific models. For instance, although Alpagsaus (Chen et al., 2024) performs excellently in inner comparison, its performance is unstable when compared against multiple external LLMs.

### v3

@@ 3.4. Methods Based on Small Models
MoDS (MoDS, 8) focuses on instruction selection through three criteria: quality (instructional data fidelity), coverage (variety of instruction types), and necessity ( impact of instruction on LLM fine-tuning). The process unfolds in four key steps. Firstly, employing a reward model to evaluate the instruction dataset quality, selecting a subset denoted as $D_{h}$, comprising instructions that surpass a predefined quality threshold. Secondly, utilizing the k-center-greedy algorithm (k-center-greedy, 23) to identify seed instructions, which ensures a diverse and representative sample of the instruction dataset. Thirdly, fine-tuning a pre-trained LLM with the seed instructions. Subsequently, this refined model is applied to $D_{h}$ to generate a new dataset, $D_{inference}$. This dataset is then evaluated using the reward model, which identifies instructions critical for learning of the LLM by focusing on those with lower scores. A threshold is established to select Augmented Instruction Data, which is specifically tailored to enhance performance of the model.Finally, the seed instructions with the Augmented Instruction Data to form a high-quality instruction subset that is poised to facilitate effective fine-tuning of the LLM.
The work (0.5data, 2) proposes a coreset-based and task-related data selection method: firstly, obtain sentence embeddings of samples by pre-trained language models (like BERT), then select the center point by applied unsupervised clustering on these embeddings, and finally use the KCenterGreedy (KCentergreedy, 24) algorithm to retrieve core samples from the given dataset. This method effectively reduces the required training data volume, maintaining or potentially enhancing model performance.
@@ 3.5. Comparative Analysis of Data Selection Methods
Model
Dependence
Model-agnostic
and heuristic
Requires supervised
model training
Depends on external
black-box APIs
(e.g., GPT-4)
Uses small pre-
trained models
(e.g., BERT, TinyLM)
@@
Domain
Adaptability
Low High Medium Low to Medium
@@
Selection
Effectiveness
Variable High High Medium
@@
Indicator-based methods clearly excel in scalability: they are fast, parallelizable, and suitable for large-scale filtering without model inference. Trainable methods, by contrast, face serious scalability challenges due to the high computational cost of model training and inference. Black-box LLM-based methods suffer from scalability bottlenecks due to API latency, usage limits, and costs, making them impractical for processing massive datasets unless sufficient resources are available. Small model methods offer a middle ground—they are not as scalable as pure heuristics but can be applied to larger datasets with reasonable efficiency.
@@
When it comes to interpretability, heuristic indicator methods clearly lead, offering intuitive, transparent criteria for data selection (e.g., length, diversity, perplexity). This is important for tasks requiring human oversight or explainable decisions. In contrast, both trainable and black-box LLM-based methods operate as opaque systems, making it difficult to trace why certain instances are chosen. Small models offer moderate interpretability—they can be analyzed via model outputs and saliency techniques but are less transparent than heuristics.
@@ 4. Evaluation methods and Result Analysis
Table 2. Performances of different selection methods on wining rate.
Selection methods Model pairs for comparison Training set Win rate across various benchmarks Vicuna/Koala/WizardLM/Self-inst/LIMA total WS IFD (IFD, 11) Llama-7B(5%), Llama-7B(full) alpaca 1.125/0.97/1.077/1/1.1 1.04 Llama-7B(10%), Llama-7B(full) 1.037/1.055/1.114/1.123/1.103 1.097 Llama-7B(15%), Llama-7B(full) 1/1.038/1.114/1.027/1.09 1.064 Llama-7B(10%), Llama-7B(full) WizardLM 1.1625/1.1278/1.1147/1.0278/1.1067 1.0971 Llama2-7B(5%), Llama2-7B(full) alpaca 1.5875/1.4889/1.4266/1.2937/1.4733 1.4311 Random Sampling Llama-7B(5%), Llama-7B(full) alpaca - 0.9 InstructionGPT4 (InstructionGPT-4, 29) miniGPT4(6%), miniGPT4(full) cc_sbu_align - 1.167 Alpagsaus (AlpaGasus, 3) Llama-7B(9k), Llama-7B(full) alpaca 1.2125/1.0222/1.0596/1.0556/- 1.0658 Llama-7B(9k), Llama-7B(3k) 1.1/1.183/1.082/1.17/- 1.074 Llama-7B(9k), Llama-7B(6k) 1.05/1.072/1.05/1.087/- 1.082 Llama-13B(9k), Llama-13B(full) 1.2125/1.0167/1.133/1.0198/- 1.074 MoDS (MoDS, 8) Llama2-7B(2k), Llama2-7B(full) alpaca 1.7125/1.5111/1.4725/1.369/1.4933 1.4786 InstructionMining (InstructionMining, 1) Llama-7B(2k), Llama-7B(full) dolly - 1.088
  • ’-’ denotes values not reported in the original paper, and model(x) denotes tuning the model on the given training set using x samples or x% of samples.
@@ 4.1. Wining Rate
To evaluate the effectiveness of the dataset selection methods, (AlpaGasus, 3) define the winning rate of the LLM fine-tuned with less samples (i.e., LLM-sub) compared to the LLM fine-tuned with full samples (i.e., base LLM) and is calculated as:
(10) $\text{Win Rate}=\frac{Num(win)-Num(lose)}{Num(all)}+1,$
@@
LLM-sub represents the LLM fine-tuned on the subset that filtered from the training set by the selection methods while the base LLM often involves 2 types: i) fine-tuned on the full training set and ii) fine-tuned on the same-scale subset filtered by the regular selections (e.g. random sampling and instruction length). The outputs of LLM-sub and the base LLM are rated by the judge on a scale from 1 to 10, and usually GPT4 is employed as the judge. To address the positional bias of the judge, IFD (IFD, 11) send the outputs of these two LLMs to the judge twice with different orders. According to this work, the wining cases refer to LLM-sub outperforming the base LLM in both times or winning in one and tying in the other. The losing cases refer to LLM-sub lagging behind the base LLM in both times or tying in one and losing in the other. The wining rates of different selection methods on the testing benchmarks are summarized in Table 2.
@@
Selection methods Training set(samples) Base model Win rate across various benchmarks total WS activeIT (ActiveIT, 10) selfinstruct(2k) Llama-7B 1.107 selfinstruct(full) 1.293 BBH/DROP/MMLU/Human-Eval/Avg DQ (DatasetQuantization, 37) alpaca(20%) Llama-7B 32.7/26.7/39.8/9.2/27.1 alpaca(full) 32.9/26.3/41.6/10/27.7 Vicuna RS/Vicuna WTR/Koala RS/ Koala WTR DIVERSEEVOL (Self-Evolved, 30) Dolly(1k) Llama-7B 79.69/20/62.29/6.67 Dolly(full) 73.84/5/57.9/3.33 SelfIns(1k) 79.16/7.5/66.95/6.11 SelfIns(full) 73.03/2.5/69.5/3.89 HellaSwag/ARC/TruthfulQA/MMLU LIFT (LIFT, 32) Platypus(15k random) 0.82/0.607/0.438/0.625 Platypus(15k) 0.844/0.643/0.49/0.645 RTE/CB/ANLI R1/ANLI R2/ANLI R3 coreset (0.5data, 2) P3(0.5) Galactica-1.3B 74.73/73.21/49.6/41.9/43.75 P3(full) 76.17/75/44/35.7/39.42
  • dataset(x) denotes tuning the base model on the x samples of the given set.
@@ 4.3. External Comparison
Methods Training set(samples) Base model MT-bench AlpacaEval
gpt-4 8.99 95.28
gpt-3.5-turbo 7.94 91.36
alpaca-13B alpaca Llama-13B 4.53 -
NUGGETS(One_Shot_Learning, 13) alpaca(7.5k) Llama-7B 5.34 -
TAGLM-13B-v1.0 (InsTag, 16) mixture(6k) Llama-13B 6.44±0.04 72.8
TAGLM-13B-v2.0 (InsTag, 16) Llama2-13B 6.55±0.02 -
Instruction Length Llama-13B 5.89 -
Random Sampling 5.84 -
IFD(IFD, 11) 5.91 -
Instruction Node(InstructionNode, 35) 5.65 -
DEITA (DEITA, 14) mixture(10k) Llama2-13B 6.79 81.09
External comparison helps reveal the generalization ability of data selection methods across different base models and architectures. For example, as shown in Table 4, IFD(IFD, 11) demonstrates robust performance when transferring from Llama-7B to Llama2-7B, indicating that the selected data subsets contain general instruction information beneficial to multiple models. However, this evaluation also exposes some limitations: certain methods that perform well in inner comparison show diminished advantages in external comparison, possibly due to overfitting to the training distribution of specific models. For instance, although Alpagsaus (AlpaGasus, 3) performs excellently in inner comparison, its performance is unstable when compared against multiple external LLMs.